### v1

# SynLogic: Synthesizing Verifiable Reasoning Data at Scale for Learning Logical Reasoning and Beyond

Abstract
Recent advances such as OpenAI-o1 and DeepSeek R1 have demonstrated the potential of Reinforcement Learning (RL) to enhance reasoning abilities in Large Language Models (LLMs). While open-source replication efforts have primarily focused on mathematical and coding domains, methods and resources for developing general reasoning capabilities remain underexplored. This gap is partly due to the challenge of collecting diverse and verifiable reasoning data suitable for RL. We hypothesize that logical reasoning is critical for developing general reasoning capabilities, as logic forms a fundamental building block of reasoning. In this work, we present SynLogic, a data synthesis framework and dataset that generates diverse logical reasoning data at scale, encompassing 35 diverse logical reasoning tasks. The SynLogic approach enables controlled synthesis of data with adjustable difficulty and quantity. Importantly, all examples can be verified by simple rules, making them ideally suited for RL with verifiable rewards. In our experiments, we validate the effectiveness of RL training on the SynLogic dataset based on 7B and 32B models. SynLogic leads to state-of-the-art logical reasoning performance among open-source datasets, surpassing DeepSeek-R1-Distill-Qwen-32B by 6 points on BBEH. Furthermore, mixing SynLogic data with mathematical and coding tasks improves the training efficiency of these domains and significantly enhances reasoning generalization. Notably, our mixed training model outperforms DeepSeek-R1-Zero-Qwen-32B across multiple benchmarks. These findings position SynLogic as a valuable resource for advancing the broader reasoning capabilities of LLMs. We open-source both the data synthesis pipeline and the SynLogic dataset at https://github.com/MiniMax-AI/SynLogic.
1 Introduction
Figure 1: The framework of logic data synthesis. The process begins with the selection of suitable tasks and the identification of key parameters that control task difficulty. Next, logic instances are generated with appropriate difficulty control (e.g., setting the grid size of Sudoku to 7). These instances are subsequently formalized into natural language instructions. Each task is paired with a task-specific verifier to check the correctness of responses. This framework enables the systematic synthesis of high-quality logic data, covering a wide range of difficulty levels and 35 task types.
The success of Deepseek R1 (DeepSeek-AI et al., 2025) and OpenAI-o1 (Jaech et al., 2024) demonstrates the great potential of post-training in advancing strong reasoning capabilities. These works reveal that the core methodology behind these advancements is reinforcement learning with verifiable rewards (RLVR), inspiring numerous replication efforts focused on RL training. However, most of these works have concentrated on the mathematics and coding domains, primarily because it is straightforward to design binary reward rules in these areas (Zeng et al., 2025b, a; Yu et al., 2025; Hu et al., 2025; Zhang et al., 2025). To foster more general and comprehensive reasoning abilities, it is essential to utilize diverse tasks and examples with verifiable rewards. In this work, we concentrate on logical reasoning as a promising domain for this objective, hypothesizing that logical reasoning serves as a fundamental building block for developing general reasoning skills. Although prior work has explored RL training in the context of logic tasks (Xie et al., 2025b; Pan et al., 2025), these efforts have typically focused on a single task, leaving the potential of broader and more diverse synthetic logic datasets largely underexplored.
Synthetic logic data presents distinct advantages and challenges. Its synthetic nature allows for unlimited data generation with controllable difficulty levels, enabling the creation of increasingly challenging samples. Additionally, the intrinsic properties of some logic tasks like Sudoku often require trial and backtracking in the reasoning process, which closely relates to the “aha moments” (DeepSeek-AI et al., 2025) in problem-solving. Therefore, the primary advantages of synthetic logic data for RLVR lie in its scalability and inherent characteristics that align well with complex reasoning processes. The main challenge, however, is the complexity of generating and designing specific rules for different logic tasks, as tasks like Game of 24 and Sudoku each require distinct verifiers.
Recent works primarily focus on logic evaluation (Suzgun et al., 2022; Ma et al., 2024; Kazemi et al., 2025), but lack high-quality accessible logical reasoning training data. In this work, to address the gap in comprehensive logic tasks, we present SynLogic: a logical reasoning data synthesis framework and a comprehensive synthetic logic dataset containing 35 tasks, including typical logical tasks such as Sudoku, Game of 24, and Cipher. For each task, we develop task-specific generation code paired with a corresponding rule-based verifier, allowing for fine-grained difficulty control through adjustable generation hyperparameters.
To validate the effectiveness of reinforcement learning on the SynLogic data, we run RL training on it with the GRPO algorithm (Shao et al., 2024) and implement binary outcome rewards determined by each task’s verification rules. By adapting recent GRPO training techniques introduced in DAPO (Yu et al., 2025), we successfully train Qwen2.5 Base models (Yang et al., 2024) on the SynLogic data in a zero RL training setting, achieving progressively longer COT responses and observing the emergence of reflection behaviors. Starting from Qwen2.5-7B-Base and Qwen2.5-32B-Base foundations, our models achieve over 8 absolute percentage points improvement on the logic benchmark KOR-Bench (Ma et al., 2024) compared to their instruction models. Notably, our 32B model surpasses DeepSeek-R1-Distill-Qwen-32B on BBEH (Kazemi et al., 2025) tasks by 5 absolute points, establishing SynLogic as the state-of-the-art open-source dataset for logical reasoning to date. Additionally, both models demonstrate strong generalization to unseen mathematics domains.
Furthermore, we explore mixing the SynLogic data with mathematics or coding data for RL training. Surprisingly, conducting the mixed training on Qwen2.5-7B-Base model (Yang et al., 2024), incorporating SynLogic data improves training efficiency for developing mathematical and coding skills. For mathematics, mixed training maintains similar mathematics performance under the same number of training steps, which consume fewer math training samples. Simultaneously, mixed training achieves much higher performance on logic tasks. A similar trend is observed when mixing SynLogic with coding data, further demonstrating the complementary benefits of logical reasoning training. Finally, we conduct large-scale mixed training on the Qwen2.5-32B-Base model to enhance the capability of Zero-RL training. Our mixed training achieves superior performance on multiple benchmarks compared to the DeepSeek-R1-Zero-Qwen-32B model, consistently outperforming or matching it on BBEH (Kazemi et al., 2025), KOR-Bench (Ma et al., 2024), LiveCodeBench (Jain et al., 2025), and GPQA-Diamond (Rein et al., 2024), validating the generalization benefits provided by the inclusion of logical reasoning data.
2 SynLogic: Synthesizing Logical Reasoning Data at Scale
Table 1: Comparison of SynLogic with existing synthetic logic datasets. *The number of tasks of KOR-Bench is based on the broader categorization in the paper. “Trainable” indicates whether the dataset provides training data.
Dataset Tasks Trainable Adjustable Difficulty
BBH (Suzgun et al., 2022) 23
Zebra Logic (Lin et al., 2024) 1
KOR-Bench (Ma et al., 2024) 5*
K&K (Xie et al., 2025a) 1
BBEH (Kazemi et al., 2025) 23
SynLogic 35
2.1 Background
Logical reasoning has long been a crucial indicator of model intelligence (ai2, 2019; Suzgun et al., 2022), valued for both its significance and synthetic accessibility. With the advancement of reasoning capabilities in Large Language Models (LLMs), researchers have developed increasingly challenging benchmarks to evaluate logical reasoning abilities (Kazemi et al., 2025; Ma et al., 2024). However, as illustrated in Table 1, existing benchmarks either lack training support or are limited to a small number of tasks. Synthetic logic data serves as an important source of verifiable data and offers straightforward control over task difficulty, presenting the potential for developing scalable stronger models by training on it. Consequently, comprehensive synthetic logic datasets are essential for developing general strong reasoning models (Seed et al., 2025).
2.2 The Data Synthesis Framework
To synthesize large-scale, diverse synthetic data, we develop a comprehensive data synthesis framework encompassing 35 tasks. While the benchmarks in Table 1 include a wide variety of tasks, a significant challenge we faced is that nearly all evaluation benchmarks do not open-source their data generation methods. This hinders us from building training data for these logic tasks directly. Therefore, we develop these tasks independently, building the SynLogic framework, illustrated in Figure 1. The framework consists of the following key components:
  1. 1.
    Task Selection We select a diverse set of logic tasks that require non-trivial reasoning, drawing from two carefully curated categories of data sources: (1) widely recognized puzzle problems from logic communities, such as the game of 24, Sudoku, and cryptarithms. Many of these puzzles have been previously highlighted in works like (Kurtic et al., 2024; Li et al., 2024; Ma et al., 2024). (2) Logic tasks featured in established evaluation benchmarks, including BBH (Suzgun et al., 2022) and BBEH (Kazemi et al., 2025). Detailed descriptions and sources for all 35 tasks can be found in the Appendix A.1.
  2. 2.
    Parameter Identification For each task, we identify key parameters that control difficulty (e.g., grid size in Sudoku, or missing numbers in Math Path). These parameters form the basis for scalable and adjustable difficulty data synthesis.
  3. 3.
    Logic Instance Generation We formalize the task-specific rules into code by manually implementing rule-based logic generators for each task. These generators are designed to encode the specific constraints and rules of the logic problems, ensuring that the generated instances adhere to the intended task structure (e.g., enforcing the unique digits rule in Sudoku). This rule-based approach allows us to efficiently produce large quantities of data and to cover a broad spectrum of difficulty levels by adjusting the difficulty related parameters. All generated instances undergo automated checks for correctness and solvability.
  4. 4.
    Appropriate Difficulty Control To ensure that the generated data is both challenging and learnable, we carefully adjust difficulty-related parameters during data generation. We use strong reasoning models, DeepSeek R1 (DeepSeek-AI et al., 2025) and OpenAI-o3-mini to set an upper bound on difficulty: the highest difficulty parameters for which R1 or o3-mini can solve samples with a pass@10 greater than zero, representing the limit of these models’ solvability. This approach prevents the inclusion of instances that are too difficult. Similarly, we use chat models to determine the lower bound of difficulty: the lowest difficulty parameters for which the models achieve a pass rate between 0 and 0.5. This dual-bound approach ensures that the dataset includes a balanced range of samples, maintaining an appropriate level of complexity and learnability.
  5. 5.
    Prompt Formalization To facilitate training and evaluation with LLMs, we convert abstract logic instances into natural language prompts using task-specific prompt templates. This step ensures that each instance is accessible to both humans and language models.
  6. 6.
    Verification Suite For every task, we implement a dedicated verifier that can automatically check the correctness of model outputs, supporting both training supervision and automatic evaluation.
A key innovation in our approach is the development of customized difficulty control mechanisms for each task type. Unlike existing benchmarks that often provide fixed-difficulty evaluation data, our system allows precise calibration of problem complexity through task-specific parameters, such as grid size in Sudoku. This difficulty-tuning capability enables the creation of different difficulty level data, presenting the potential of progressively challenging training curricula. We overcame significant challenges in implementing these controls, as many evaluation benchmarks do not open-source their data generation methods. At last, most tasks in SynLogic are designed with: (1) a data generation code capable of producing varied instances, (2) a corresponding verification rule for evaluating solution correctness, and (3) configurable difficulty parameters to enable controlled difficulty of generated data. We independently develop and generate data for 33 tasks in our dataset, while only the data of 2 tasks (Zebra Puzzle (Lin et al., 2024) and ARC-AGI (Chollet, 2019)) are directly adopted from existing open source resources.
Risk of Data Contamination
Although several tasks overlap between our selected tasks and current benchmarks, such as KOR-Bench and BBEH, the synthetic nature of our data, combined with the large synthesis space, makes the probability of generating data identical to benchmark test samples very low – we have verified that there are no identical samples between our generated datasets and the benchmark test sets.
2.3 The SynLogic Datasets
We synthesized our dataset with controlled difficulty parameters for each task, carefully balancing challenge and learnability to ensure the success of our subsequent experiments §3. To accommodate different model capacities, we developed two distinct versions of our dataset: SynLogic-Hard for Qwen2.5-32B training and SynLogic-Easy for Qwen2.5-7B training. SynLogic-Hard presents more complex challenges with its broader difficulty level for each tasks with difficulty upper bound described in § 2.2. For SynLogic-Easy, we systematically lower difficulty parameters across all tasks to create easier versions. Despite these adjustments, eight tasks still remain beyond the learning capacity of the 7B model with zero training accuracy after RL, thus we removed them from this easy version, where the details about the removed tasks are provided in Appendix A.2. Finally, we synthesized 33k SynLogic-Hard samples and 16k SynLogic-Easy samples used in subsequent experiments for training, along with 10 validation samples per task, separately for the Easy and Hard validation splits.
Difficulty Analysis
To assess the difficulty of the synthetic data, we conduct an evaluation on the validation splits, assessing model performance using both avg@8 (average pass rate with eight attempts) and pass@8 (success within eight attempts) metrics. The results, illustrated in Figure  2, confirm the appropriate difficulty levels for each model scale, demonstrating that our datasets provide suitable training challenges across different model capacities.
(a) 7B models on SynLogic-Easy
(b) 32B models on SynLogic-Hard
Figure 2: Evaluation of task difficulty across our dataset versions. (a) Shows the performance of 7B-scale models on the SynLogic-Easy dataset, while (b) demonstrates the performance of 32B-scale models on the more challenging SynLogic-Hard dataset. Results are measured using avg@8 (average pass rate with eight attempts) and pass@8 (success within eight attempts) metrics, illustrating the appropriate difficulty control for each model scale.
3 Reinforcement Learning on SynLogic
Table 2: Evaluation accuracy (%) of 7B and 32B models across logic benchmarks (SynLogic-Val, KOR-Bench, BBH, BBEH) and mathematical benchmarks (AIME 2024, MATH 500, AMC 2023). For 7B models, SynLogic-Val refers to SynLogic-Easy, while for 32B models, it refers to SynLogic-Hard as described in § 2.3. All evaluations were conducted under zero-shot conditions, with SynLogic-Val and AIME 2024 reported as avg@8 to reduce variance.
Model Logic Benchmarks Mathematical Benchmarks
SynLogic-Val KOR-Bench BBH BBEH AIME 2024 MATH 500 AMC 2023
Qwen2.5-7B-Base 2.8 11.6 45.2 3.8 0.3 64.6 30.0
Qwen2.5-7B-Instruct 9.0 38.6 62.7 12.4 6.3 76.4 52.5
SynLogic-7B 44.4 48.1 66.5 8.0 10.0 71.8 55.0
Qwen2.5-32B-Base 1.6 10.9 58.4 3.3 4.5 68.6 45.0
Qwen2.5-32B-Instruct 12.0 54.7 84.5 17.5 10.0 82.2 57.5
R1-Distill-Qwen-32B 33.0 66.6 88.3 19.2 72.6 94.3 85.0
SynLogic-32B 52.9 62.2 85.8 25.5 19.6 82.0 57.5
Reinforcement learning with verifiable rewards (RLVR) has emerged as a highly effective approach for enhancing reasoning capabilities in large language models (Zeng et al., 2025b; Yu et al., 2025; Hu et al., 2025; Zhang et al., 2025). Building on these advances, our experimental framework also focuses on applying reinforcement learning techniques to the SynLogic dataset, leveraging the verifiable nature of logical reasoning tasks. In this section, we validate the effectiveness of reinforcement learning training on the SynLogic dataset using Qwen2.5-7B-Base and Qwen2.5-32B-Base models.
3.1 Setup Details
Training Template
Following the DAPO training prompt (Yu et al., 2025), we modify and design the training prompt template for logic training as shown in Figure 3:
[…]
Figure 3: The prompt template used for training models on SynLogic data.
Reward Design
Our reward function employs a binary scoring mechanism that evaluates both format adherence and answer correctness. Specifically, we assign a reward of 1 only when a model-generated response satisfies two criteria: (1) it correctly follows the designated format by including both <think> </think> and <answer> </answer> tags, and (2) the final answer provided is correct. Responses that either deviate from the required format or contain incorrect answers receive a reward of 0.
$R(o_{i})=\begin{cases}1,&\text{if }\text{format}(o_{i})=\text{True }\wedge\text{ correct}(o_{i})=\text{True}\\ 0,&\text{otherwise}\end{cases}$ (1)
where $\text{format}(o_{i})$ evaluates whether response $o_{i}$ includes both the required <think> </think> and <answer> </answer> tags, and $\text{correct}(o_{i})$ determines whether the answer provided is correct verified by its task’s verification rule.
Training Details
For our experiments, we synthesized approximately 16k SynLogic-Easy and 33k SynLogic-Hard instances to train the Qwen2.5-7B-Base and Qwen2.5-32B-Base models with DAPO, respectively, as described in §2.3. During training, we employed a prompt batch size of 128, generated 16 rollouts per prompt, and set maximum rollout lengths of 16,384 tokens for the 7B model and 28,672 tokens for the 32B model. We configured the clip high parameter $\epsilon_{\text{high}}$ at 0.28. Additional training hyperparameters and implementation details are provided in Appendix B.1.2.
Evaluation Details
Our evaluation strategy encompasses two distinct benchmark categories. For assessing logical reasoning capabilities, we employ the validation splits of SynLogic alongside established benchmarks including Knowledge-Orthogonal Reasoning (KOR-Bench) (Ma et al., 2024), BBH (Suzgun et al., 2022), and the substantially more challenging BBEH (Kazemi et al., 2025). To investigate cross-domain generalization effects, we incorporate mathematics evaluations on MATH 500 (Hendrycks et al., 2021), AMC 2023, and AIME 2024. All evaluations are conducted in a zero-shot setting, with avg@8 metrics computed for AIME 2024 and SynLogic-Val to mitigate variance.
3.2 Results
(a) Avg Length and Reflection of 7B Training.
(b) Avg Length and Reflection of 32B Training.
Figure 4: Response length and reflection ratio across the 7B and 32B training process on the training dataset. The reflection ratio represents the proportion of generated responses containing at least one reflection phrase (including “recheck”, “rethink”, “try again”, “let’s correct it”, “re-evaluate”, “check again”, “think again”).
Substantial Improvements in Logical Reasoning
The evaluation results presented in Table 2 demonstrate significant improvements across logical reasoning tasks. Beyond the notable gains on SynLogic’s validation split, our models demonstrate enhanced performance across multiple logical benchmarks, leading state-of-the-art results among open-source datasets. Our 7B model achieves 48.1% on KOR-Bench (Ma et al., 2024), outperforming Qwen2.5-7B-Instruct by nearly 10 absolute percentage points. Similarly, our 32B model surpasses Qwen2.5-32B-Instruct by 7 percentage points on KOR-Bench. Notably, our 32B model exceeds R1-Distill-Qwen32B (DeepSeek-AI et al., 2025) by 6 percentage points on the challenging BBEH benchmark (Kazemi et al., 2025), showcasing the effectiveness of the SynLogic dataset in driving state-of-the-art logical reasoning performance.
Generalization to Mathematical Domains
Our experimental results demonstrate significant generalization capabilities to mathematical domains, as shown in Table 2. Despite being primarily trained for logical reasoning, SynLogic models exhibit strong performance across mathematical benchmarks over their base models. SynLogic-7B achieves 10.0% on AIME 2024, a nearly 10 absolute point improvement compared to Qwen2.5-7B-Base (0.3%), 71.8% on MATH 500, a 7.2 absolute point gain, and 55.0% on AMC 2023, a 25-point increase. More remarkably, SynLogic-32B achieves 19.6% on AIME 2024, a 4.4x improvement over Qwen2.5-32B-Base (4.5%), while its performance on MATH 500 (82.0%) and AMC 2023 (57.5%) shows substantial gains of 13.4 and 12.5 points, respectively. Without mathematics training data, our models nearly match or surpass the instruction models, suggesting that enhancements in logical reasoning capabilities transfer effectively to mathematical problem-solving. This aligns with the observation in Logic-RL (Xie et al., 2025b), highlighting the fundamental connection between logical and mathematical reasoning skills.
Increased Chain-of-Thought Length
As shown in Figure 4, recording the response length and reflection ratio during the training process reveals that training on SynLogic data leads to stable increases in response length for both models. The 7B model reaches an average of approximately 2500 tokens, while the 32B model achieves around 4000 tokens. Additionally, the increasing reflection ratios also indicate the emergence of cognitive behaviors during training (Gandhi et al., 2025). Both the extended response lengths and increased prevalence of reflection tokens suggest that synthetic logic reasoning tasks inherently align with the long-thinking paradigm of LLM. Detailed training accuracy results are provided in Appendix B.1.3.
4 Scaling RL Training with Diverse Verifiable Reasoning Data
Having verified the success of reinforcement training on SynLogic alone, we now leverage verifiable reasoning data from math, coding, and logical reasoning domains, scaling RL training with diverse verifiable reasoning data. Concretely, we mix SynLogic with mathematical or/and coding training data, and perform RL training on the mixed datasets. We will study how combining logical reasoning with code and mathematical data influences training efficiency on 7B models and enhances the Zero-RL capabilities for 32B models.
4.1 Setup Details
For mathematical training data, we directly utilize the 17k samples provided in DAPO (Yu et al., 2025). For coding data, we assembled approximately 9k samples from various online coding platforms such as Codeforces. We adapted a similar prompt template to that used in our SynLogic training; the detailed templates are presented in Appendix B.1. We maintained the same reward design approach as described in  §3.1. Specifically, for coding tasks, the reward is set as 1 if the output is correctly formatted and all test cases pass; otherwise, the reward is 0.
(a) Acc on KOR-Bench.
(b) Avg Acc on Math.
(c) Avg Math Acc vs. Consumed Math Data
Figure 5: Performance comparison of 7B models trained on mixed data (Logic+Math) versus math-only (Math) data. (a) Accuracy on KOR-Bench. (b) Average accuracy across three mathematics benchmarks (MATH 500, AIME 2024, AMC 2023) as a function of training steps. (c) Average accuracy on mathematics benchmarks as a function of consumed mathematical data volume.
(a) Acc on KOR-Bench.
(b) Avg Acc on Code.
(c) Avg Code Acc vs. Consumed Coding Data.
Figure 6: Performance comparison of 7B models trained on mixed data (Logic+Coding) versus coding-only (Coding) data. (a) Accuracy on KOR-Bench. (b) Average accuracy across two coding benchmarks (validation split of our coding data and LiveCodeBench (Jain et al., 2025)) as a function of training steps. (c) Average accuracy on coding benchmarks as a function of consumed coding data volume.
4.2 Mixing SynLogic with Math or Code Data: A Pilot Ablation Study
Mixed Training with Math
We sample approximately 17k samples from SynLogic-Easy and combine them with 17k math data for training on the Qwen2.5-7B-Base model. For controlled comparison, we also conduct reinforcement learning using exclusively math data. Both experimental configurations maintain identical hyperparameters, optimization settings, and computational resources to ensure fair evaluation. Figure 5 presents a comparison of training dynamics. Running for the same number of training steps, mixed training (Logic+Math) achieves comparable performance to math-only training on average across three mathematical benchmarks (Figure 5b), while consuming fewer math samples. Under the same volume of processed math data, mixed training achieves higher accuracy (Figure 5c). Moreover, mixed training steadily improves logical reasoning, as reflected in rising KOR-Bench scores (Figure 5a), which are nearly 10 absolute percentage points higher than those achieved with math-only training. These results suggest that mixed training facilitates more efficient optimization, potentially due to shared abstract reasoning mechanisms across domains.
Mixed Training with Code
Following a similar methodology, we sample approximately 9k samples from SynLogic-Easy and combine them with 9K code samples to train the Qwen2.5-7B-Base model. As a control, we conduct parallel training using exclusively coding data. Both training configurations maintain identical parameters to ensure a fair comparison. To measure the coding ability, we include the validation split of our coding data and the LiveCodeBench (Jain et al., 2025) (the same version as used in DeepSeek’s report (DeepSeek-AI et al., 2025)) for evaluation. As shown in Figure 6, we observe a similar phenomenon of more efficient training dynamics when mixing code with SynLogic-Easy. Models trained on Logic+Coding data achieve higher performance on coding benchmarks than code-only training when consuming the same volume of coding data. Simultaneously, mixed training improves logical reasoning, as evidenced by 10 absolute points better KOR-Bench scores (Figure 6a). These findings reinforce the complementary nature of logical reasoning in enhancing domain-specific capabilities.
4.3 32B Zero-RL Training with Diverse Reasoning Data
Table 3: Performance comparison across multiple benchmarks. The evaluation metrics vary by dataset: BBEH (Kazemi et al., 2025) uses pass@1, while KOR-Bench (Ma et al., 2024), LiveCodeBench (LCB)(Jain et al., 2025), and GPQA-Diamond(Rein et al., 2024) use avg@4. AIME 2024 is evaluated using avg@8. All training configurations (Zero-Mix-2 and Zero-Mix-3) are run for the same number of training steps to ensure a fair comparison of results.
Model BBEH KOR-Bench LCB AIME 2024 GPQA Diamond
DeepSeek-R1-Distill-Qwen-32B 19.2 66.6 57.2 72.6 63.1
DeepSeek-R1-Zero-Qwen-32B - - 40.2 47.0 55.0
Zero-Mix-2 (Math+Coding) 18.5 58.6 39.5 34.5 55.2
Zero-Mix-3 (SynLogic+Math+Coding) 28.6 65.0 40.7 35.8 57.5
Building on the previous observation, here we scale up the diverse, verifiable training data by mixing math, coding, and SynLogic datasets, and perform RL training on the Qwen2.5-32B-Base model. Specifically, we use a mix of 35k mathematical samples, 9k coding samples, and 17k SynLogic samples for training. We term this training configuration as Zero-Mix-3. We additionally conduct a Zero-Mix-2 setting that only mixes coding and mathematical data, serving as an ablation baseline to study the effect of SynLogic in such a scalable setting. Related to our Zero-Mix-2 setup,  Zhang et al. (2025) recently demonstrated that combining mathematical data with coding data is able to facilitate coding learning. Here we further scale up this trend by including our proposed SynLogic dataset. To evaluate generalization, we include an out-of-domain benchmark, GPQA Diamond (Rein et al., 2024), to study how the addition of SynLogic impacts broader reasoning capabilities. Both Zero-Mix-3 and Zero-Mix-2 configurations are run for the same number of training steps to ensure a fair and controlled comparison.
Results
As shown in Table 3, the Zero-RL training in Zero-Mix-3 (SynLogic+Math+Coding) achieves superior performance across multiple evaluations. On logic benchmarks, Zero-Mix-3 nearly matches the performance of DeepSeek-R1-Distill-Qwen-32B on KOR-Bench and surpasses it by 8 points on BBEH. Notably, Zero-Mix-3 also matches DeepSeek-R1-Zero-Qwen-32B on the coding benchmark LiveCodeBench (Jain et al., 2025) and outperforms it on GPQA-Diamond. Compared to the Zero-Mix-2 (Math+Coding) experiment, Zero-Mix-3 consistently delivers higher performance across all benchmarks. Specifically, Zero-Mix-3 shows a significant improvement of over 10 points on BBEH, 6 points on KOR-Bench, and over 2 points on the out-of-domain benchmark GPQA Diamond. These results strongly validate the significant generalization benefits provided by the inclusion of SynLogic.
5 Conclusion
We present SynLogic: a data synthesis framework and a comprehensive synthetic logic dataset with 35 diverse tasks, addressing the lack of high-quality logic training data. Using SynLogic, we trained Qwen2.5 models with the GRPO algorithm, achieving significant gains on logic benchmarks like KOR-Bench and strong generalization to unseen mathematical tasks. Notably, our 32B model outperformed DeepSeek-R1-Distill-Qwen-32B on BBEH. Mixed training with SynLogic further improved training efficiency and performance, showcasing the complementary benefits of logical reasoning across domains. We hope SynLogic inspires broader exploration of synthetic datasets and logical reasoning to develop stronger reasoning capability models.
Limitations
First, due to limited computational resources, we do not conduct precise difficulty tuning for every task, meaning the curated data may not be fully optimized for training 7B and 32B models. Second, we do not implement training with dynamically adjustable difficulty, where sample complexity progressively increases during training. We hypothesize that such an approach could further enhance the logical reasoning capabilities of LLMs. Both of these directions are left as opportunities for future exploration.
Appendix A Comprehensive Overview of SynLogic
A.1 Task Composition and Sources
Table 4 presents the diverse collection of tasks incorporated in SynLogic. We have carefully selected these tasks from established benchmarks including KOR-Bench [Ma et al., 2024], BBH [Suzgun et al., 2022], and BBEH [Kazemi et al., 2025]. Additionally, we integrated some logical reasoning tasks not previously featured in these benchmarks, such as Mathador [Kurtic et al., 2024] and Minesweeper [Li et al., 2024].
Our collection comprises 35 distinct tasks, with only two (Zebra Puzzle [Lin et al., 2024] and ARC-AGI [Chollet, 2019]) using existing data sources. For all remaining tasks, we generated custom datasets. Importantly, we developed and implemented verifiers for all tasks in the collection, ensuring consistent evaluation across the benchmark.
Table 4: Tasks in the Dataset Collection with Descriptions
No. Task Name Description
1 ARC-AGI A collection of general intelligence tasks requiring abstract reasoning and pattern recognition.
2 Arrow Maze A maze-solving task where arrows dictate movement, requiring pathfinding logic.
3 Boolean Expressions Evaluating logical expressions with AND, OR, NOT operators.
4 Buggy Tables Correcting flawed data in tables based on logical constraints.
5 Calcudoko A math-based Sudoku variant with arithmetic constraints.
6 Campsite Placing tents in a grid while satisfying adjacency rules.
7 Cipher Decoding encrypted messages based on given rules.
8 Cryptarithm Solving arithmetic puzzles where letters represent digits.
9 Dyck Language Validating bracket sequences for correct nesting.
10 Dyck Language Errors Identifying and correcting errors in bracket sequences.
11 Dyck Language Reasoning Errors Advanced reasoning for errors in bracket nesting logic.
12 Futoshiki A grid-based logic puzzle with inequality constraints.
13 Goods Exchange Tracking item exchanges among multiple participants.
14 Kukurasu A grid-based puzzle involving row/column weight sums.
15 Mathador A math strategy game involving arithmetic operations.
16 Math Path Finding correct numbers to satisfy equations in a grid.
17 Minesweeper Logical deduction to uncover mines on a grid.
18 Norinori Placing domino tiles in a grid with adjacency rules.
19 Number Wall Constructing walls to separate grid regions based on rules.
20 Numbrix Filling a grid with consecutive numbers in order.
21 Object Counting Counting specific objects under given constraints.
22 Object Properties Inferring and reasoning about object attributes.
23 Operation Solving puzzles with custom-defined mathematical operations.
24 Skyscraper Puzzle Determining building heights based on visibility clues.
25 Space Reasoning Reasoning about spatial relationships in a grid.
26 Space Reasoning Tree Advanced spatial reasoning tasks with hierarchical relationships.
27 Star Placement Puzzle Placing stars in a grid while avoiding adjacency conflicts.
28 Sudoku Solving the classic number-placement puzzle.
29 Survo Filling grids to satisfy row and column sum constraints.
30 Time Sequence Scheduling tasks or events with overlapping constraints.
31 Web of Lies Determining truth-tellers and liars through logical statements.
32 Word Sorting Sorting words based on custom rules or constraints.
33 Word Sorting Mistake Identifying mistakes in word sorting logic or reasoning.
34 Wordscapes A crossword puzzle where players fill words from lists while matching intersections.
35 Zebra Puzzle Solving complex logic puzzles with multiple constraints.
A.2 SynLogic-Hard and SynLogic-Easy
The SynLogic-Hard dataset encompasses all 35 tasks, representing a challenging upper bound calibrated to the solvability thresholds of DeepSeek R1 and OpenAI-o3-mini. During our experiments with Qwen2.5-32B-Base, we observed consistent training accuracy gains across this comprehensive task set. However, this difficulty level proved excessive for smaller-scale models like Qwen2.5-7B-Base. Despite reducing the difficulty parameters across tasks, eight specific tasks (Arrow Maze, Goods Exchange, Kukurasu, Minesweeper, Norinori, Object Counting, Space Reasoning Tree, and Wordscapes) persistently yielded zero accuracy when training the 7B models. Consequently, we developed SynLogic-Easy, a modified variant that excludes these particularly challenging tasks, to provide a more appropriate training dataset for the Qwen2.5-7B-Base model.
Appendix B Training and Evaluation Details
B.1 Training
B.1.1 Training Template
We provide the training template for math and coding here in Figure 7 and Figure 8.
[…]
Figure 7: The prompt template used for training models on math data.
[…]
Figure 8: The prompt template used for training models on coding data.
B.1.2 Training hyper-parameters
For 7B model training, we adopted the following hyper-parameters: with learning rate 1e-6, GRPO group size 16, max prompt length 2048, max response length 16384, prompt batch size 128, mini batch size 64, with clip high 0.28, clip low 0.2.
For 32B model training in §3, we adopted the following hyper-parameters: with learning rate 2e-6, GRPO group size 16, max prompt length 2048, max response length 28672, prompt batch size 128, mini batch size 16, with clip high 0.28, clip low 0.2.
For 32B model training in §4, we adopted the following hyper-parameters: with learning rate 2e-6, GRPO group size 16, max prompt length 2048, max response length 12288, prompt batch size 512, mini batch size 16, with clip high 0.28, clip low 0.2.
B.1.3 Training Dynamics Analysis
We present the training accuracy progression of our 7B and 32B models on the SynLogic dataset in Figure 9. The results demonstrate a consistent improvement in accuracy throughout the training process for both model sizes.
(a) Training accuracy of 7B model
(b) Training accuracy of 32B model
Figure 9: Training accuracy progression for both model sizes on the SynLogic dataset. Both models exhibit steady improvement in performance as training progresses.
B.2 Evaluation
(a) Performance on MATH 500.
(b) Performance on AIME 2024.
(c) Performance on AMC 2023.
Figure 10: Comparative accuracy (%) of models trained with mixed data and math-only data across three mathematical benchmarks: MATH 500, AIME 2024, and AMC 2023. All evaluations of the figure use avg@8 scoring.
B.2.1 Performance Analysis of Mixed Training with Math
We analyze the training dynamics when combining our logical reasoning dataset with mathematical content. Figure 10 presents performance results across three mathematical benchmarks: MATH 500, AIME 2024, and AMC 2023. These results demonstrate that mixed training can achieve similar performance with the same number of training steps while using less mathematical data. This suggests that mixed training creates more efficient training dynamics.
B.3 Performance Analysis of Mixed Training with Coding
We also provide the training dynamics when mixed training with coding data in Figure 11. The observation is similar to mixed training with math, suggesting that mixed training leads to more efficient training dynamics.
(a) Performance on our coding data validation split.
(b) Performance on LiveCodeBench.
Figure 11: Comparative accuracy (%) of models trained with mixed data and coding-only data across two coding benchmarks: our coding data validation split and LiveCodeBench. All evaluations of the figure use avg@8 scoring.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Our claims match our experimental results.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: In the conclusion section.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: Empirical paper
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We will open-source our dataset and code for the data synthesis pipeline.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: The code, data and models are well documented and the documentation will be made publicly available alongside the assets following the review period.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We provide in the appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: Too costly to conduct multi-runs.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Details in the Appendix.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We conducted the paper confrom.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [No]
  49. Justification: The paper is about logical reasoning, no societal impact.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our dataset is not related to this.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Yes, they are properly credited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We describe details in Appendix.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [No]
  69. Justification: Not this type paper.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [No]
  74. Justification: Not required for our paper.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: Core method are developed by ourselves.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v2

@@
 All evaluations of the figure use avg@8 scoring.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: Our claims match our experimental results.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: In the conclusion section.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory assumptions and proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [N/A]
-  14. Justification: Empirical paper
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental result reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19. Justification: We will open-source our dataset and code for the data synthesis pipeline.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [No]
-  24. Justification: The code, data and models are well documented and the documentation will be made publicly available alongside the assets following the review period.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental setting/details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: We provide in the appendix.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment statistical significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [No]
-  34. Justification: Too costly to conduct multi-runs.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments compute resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: Details in the Appendix.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code of ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: We conducted the paper confrom.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [No]
-  49. Justification: The paper is about logical reasoning, no societal impact.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54. Justification: Our dataset is not related to this.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: Yes, they are properly credited.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [Yes]
-  64. Justification: We describe details in Appendix.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and research with human subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [No]
-  69. Justification: Not this type paper.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [No]
-  74. Justification: Not required for our paper.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
-  76. 16.
-    Declaration of LLM usage
-  77.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-  78. Answer: [N/A]
-  79. Justification: Core method are developed by ourselves.
-  80. Guidelines:
-    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
-    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.